OpenStudy (anonymous):

What are the amplitude, period, phase shift, and midline of f(x) = –3 sin(4x – π) + 2?

OpenStudy (anonymous):

What are your available answers? Or are there any?

OpenStudy (anonymous):

u shouldn't need it if u know how to do it @stratemeyerjr

OpenStudy (anonymous):

\[f(x) = Asin B(x-C)+D\] factor out 4 to get in that format \[f(x) = -3 \sin 4(x - \pi/4)+2\] |A| = amplitude midline: y = D Period (2Ï€)/B phase shift = C

OpenStudy (anonymous):

so amplitude is never negatvie right

OpenStudy (anonymous):

right, it's a measure of distance

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you're welcome